# Percept, Chat, and then Adapt: Multimodal Knowledge Transfer of Foundation Models for Open-World Video Recognition

### v1

@@ 4 Experiments
  R(2+1)Dr2p1d 74.10%
CLIP-ViTclip 74.70%
Timesformertimesformer 77.40%
UniFormeruniformer 77.80%
CLIP-3Dclip 78.80%
EVLfrozenclip 79.33%
ST-Adapterstadapter 79.60%
@@
In QV-Pipe, the learning rate, weight decay, batch size, warmup epochs, and total training epochs, are set to 1e-5, 0.2, 72, 5, and 120 respectively. We uniformly sample 8 frames with the resolution of 224$\times$224. During training, multi-scale crop, color jitter, and random grayscale augmentation are applied. During testing, the center crop is used. The dropout rate for the classification head is 0.5. MAP (Mean Average Precision) is the evaluation metric.
4.1 Comparison to the State of the Art
TinyVIRAT. We compared the current mainstream CNNr2p1d , Transformer modelsclip ; timesformer and popular adaptation methodsstadapter ; frozenclip on TinyVIRATtinyvirat as shown in Table 1. Among them, EVLfrozenclip and ST-Adapterstadapter are representative video adapter models. We do not apply PCA on adapter methods because these models have been equipped with adapter modules. It is observed that adaptation methods outperform other baselines. After incorporating the PCA method, improved results were achieved on R(2+1)D, UniFormer, Timesformer, CLIP-ViT, and CLIP-3D models compared to their respective baselines. Notably, the CLIP-ViT method, after PCA is incorporated, outperforms other adapter methods based on the CLIP model. The PCACLIP-3D model achieves the best results of 82.28%.
ARID. We compared our approach with previous mainstream models in the ARID dataset in Table 2. Due to the limited number of videos in ARID, which is less than 4k, using a transformer model would lead to overfittingvit . However, convolution neural networks are more suitable for fitting small datasets. Therefore, models with transformer components performed lower than models with CNN components. The performance of the CLIP-based adapter model was also lower. Among the baseline models, the R(2+1)Dr2p1d model achieved the best results. After incorporating the PCA method, the performance of all models surpassed their corresponding baseline models. With the help of text and visual knowledge, PCAR(2+1)D model achieves a Top-1 Acc of 98.9% surpassing the second-best method by +4.86%.
Table 2: Comparison with the state-of-the-art on ARIDarid
@@ 4.1 Comparison to the State of the Art
  ST-Adapterstadapter 72.67% 97.51%
CLIP-ViTclip 77.96% 97.08%
Timesformertimesformer 78.53% 96.02%
CLIP-3Dclip 78.79% 98.44%
EVLfrozenclip 79.01% 98.44%
UniFormeruniformer 86.60% 96.46%
DarkLight-ResNeXtdarklight 87.27% 99.47%
R(2+1)Ddarklight 90.45% 98.11%
R(2+1)D-Lightdarklight 91.55% 99.51%
R(2+1)D-DarkLight-SAdarklight 94.04% 99.87%
@@
Table 3: Comparison with the state-of-the-art on QV-Pipevideopipe
  Model mAP
  R(2+1)Dr2p1d 25.67%
UniFormeruniformer 50.72%
CLIP-3Dclip 55.80%
Timesformertimesformer 62.84%
EVLfrozenclip 63.13%
CLIP-ViTclip 63.40%
ST-Adapterstadapter 63.42%
PCAR(2+1)D 26.30%
PCAUniFormer 55.30%
PCACLIP-3D 57.14%
PCATimesformer 62.98%
PCACLIP-ViT 65.42%
QV-Pipe. The results of QV-Pipe dataset are reported in Table 3. We can see that our proposed PCACLIP-ViT methods achieve SOTA performance compared to previous methods. It outperforms the ST-Adapter method by +2.0%. Additionally, we also showcase the results of other baseline methodsuniformer ; clip ; timesformer ; r2p1d with PCA pipeline. QV-Pipe is a complex multilabel dataset requiring global information for perceptionvideopipe . The pure convolution network can not generalize well, because convolution has limited capabilities to integrate global informationuniformer . Therefore, the performance of R(2+1)Dr2p1d is not as good as transformer-based models with global attention to QV-Pipe. By incorporating the PCA method, our model outperformed the respective baseline models, achieving state-of-the-art performance.
@@ 4.2 Ablation Studies
UniFormeruniformer $\times$ $\times$ 77.80% 86.60% 96.46% 55.30%
@@
Timesformertimesformer $\times$ $\times$ 77.40% 78.53% 96.02% 60.55%
@@
CLIP-ViTclip $\times$ $\times$ 74.70% 77.96% 97.08% 63.40%
@@
CLIP-3Dclip $\times$ $\times$ 78.80% 78.79% 97.44% 55.80%
@@
Multimodal Ablation for PCA: According to Table 4, all models have gained performance improvement when incorporating only single-modal knowledge(either visual or text), compared to their corresponding baselines. This indicates that external knowledge, whether in the form of text or visual modalities, has a positive impact on improving the accuracy of the model. Moreover, the integration of both visual and textual knowledge further boosts the models’ performance, showing the effectiveness of our proposed method.
@@
The Number of Inserted Adapt Blocks: To validate the ability of the adapt module to incorporate intermediate-level features into the backbone network, we extracted the intermediate layers of the visual feature extraction models and integrated them into the backbone. Specifically, the block num represents the total number of adapt modules inserted as shown in Fig.4. For the CLIP-3D, Timesformer, and CLIP-ViT models, we extracted the Residual Attention Blocksvit features from the middle layers. For the UniFormer model, we extracted middle layer features from SABlock layersuniformer . For the R(2+1)D model, we extracted intermediate features from BasicBlocksr2p1d . We only introduced textual knowledge in the last block of the model. The results shown in Table 7, Block Num 3 achieves the best result on three datasets. With the increased number of inserted blocks, the results improve.
Results of Visual Enhancement: The first row refers to training the baseline model using the source data. The second row represents training the baseline model using augmented visual enhanced data. The final row demonstrates the effect of incorporating multimodal knowledge into PCA. As shown in Table 8, training the baseline models using enhanced data yields higher results compared to training with the source data alone. Introducing visual single-modal knowledge in PCA further enhances performance, since the PCA visual incorporates information from both the source data and the visual enhanced data. Finally, incorporating knowledge allows the network to achieve superior results.
@@ 4.3 Visualization
We visualize the attention regions of the model using Grad-CAMcam as shown in Fig.6. After incorporating visual and textual knowledge, the model gains the ability to correct the wrong answers provided by the baseline model.
Figure 6: We use CAMcam to visualize the PCA pipeline. For each case, the top left corner of each example shows four frames sampled from the original video, while the bottom left corner shows the features obtained without incorporating external knowledge. The images in the middle represents enhanced video frames. The text box contains the textual knowledge obtained by inputting the original video into the VideoChat foundation model. The upper right corner displays features that only include external visual knowledge, while the bottom right corner shows features that only include external textual knowledge. The middle section represents features obtained by incorporating both visual and textual knowledge.
(A) is from TinyVIRAT dataset. When there is no external knowledge, the model predicts the result as ”Opening”. After incorporating visual knowledge alone, the model also misjudges the video. We introduce external textual knowledge, which directly states that the person is standing behind the car. As a result, the model can predict the correct category. When the external knowledge includes both visual and textual modalities, the focus of the feature maps shifts to the body and feet of the standing person and obtaining a higher confidence score.
(B) is from the ARID dataset. Without the introduction of external knowledge, the model struggles to perceive the actions of a person in darkness, resulting in incorrect judgments and nonintuitive feature maps. The enhanced video restores hand movements in a dark environment. Textual knowledge mentions the person using a phone, reading a book, or typing, indicating hand movements. With the incorporation of single-modal knowledge, the model still fails to determine the correct category, and the feature maps remain nonintuitive. After incorporating multimodal knowledge, the model can infer comprehensively that the person is picking something. Consequently, the model correctly infers the category. Furthermore, the feature maps intuitively focus on the position of the hand.
(C) is from the QV-Pipe dataset. Without external knowledge and with only single-modal knowledge, the model cannot predict all class labels, and the feature maps are not intuitive. The mask from segment anythingsegmentanything covers both the pipeline walls and the water flow. Textual information indicates that this is a sewer with water and some dirt sediment. The model successfully obtains the correct result under the guidance of the semantic segmentation mask and the textual information. The attention map focused on the surface of the sewer and the walls of the pipe.
@@ 5 Conclusion
In this work, we propose a general knowledge transfer framework PCA, which utilizes a series of foundation models as knowledge containers and progressively incorporates and integrates external multimodal knowledge to enhance the performance of open-world video recognition. First, in the Percept stage, we enhance open-world videos to reduce domain gaps and extract external visual knowledge. Second, in the Chat stage, we generate rich semantic information about open-world videos. These diverse texts serve as external textual knowledge. Finally, in the Adapt stage, we flexibly insert multimodal knowledge adaptation modules into the backbone networks to fuse external multimodal knowledge for open-world recognition. To validate the effectiveness of our PCA framework, we conduct extensive experiments on three challenging video benchmark datasets in open-world scenarios, including the low-resolution TinyVIRATtinyvirat , the low-light environment ARIDarid , and the unconventional domain QV-Pipevideopipe . Experimental results demonstrate that our approach achieves state-of-the-art performance on these three datasets.

### v2

@@ 4.1 Comparison to the State of the Art
R(2+1)D[r2p1d] 74.10%
CLIP-ViT[clip] 74.70%
Timesformer[timesformer] 77.40%
UniFormer[uniformer] 77.80%
CLIP-3D[clip] 78.80%
EVL[frozenclip] 79.33%
ST-Adapter[stadapter] 79.60%
@@
Table 2: Comparison with the state-of-the-art on QV-Pipe[videopipe]
Model mAP
R(2+1)D[r2p1d] 25.67%
UniFormer[uniformer] 50.72%
CLIP-3D[clip] 55.80%
Timesformer[timesformer] 62.84%
EVL[frozenclip] 63.13%
CLIP-ViT[clip] 63.40%
ST-Adapter[stadapter] 63.42%
PCAR(2+1)D 26.30%
PCAUniFormer 55.30%
PCACLIP-3D 57.14%
PCATimesformer 62.98%
PCACLIP-ViT 65.42%
TinyVIRAT. We compared the current mainstream CNN[r2p1d], Transformer models[clip, timesformer] and popular adaptation methods[stadapter, frozenclip] on TinyVIRAT[tinyvirat] as shown in Table 2. Among them, EVL[frozenclip] and ST-Adapter[stadapter] are representative video adapter models. We do not apply PCA on adapter methods because these models have been equipped with adapter modules. It is observed that adaptation methods outperform other baselines. After incorporating the PCA method, improved results were achieved on R(2+1)D, UniFormer, Timesformer, CLIP-ViT, and CLIP-3D models compared to their respective baselines. Notably, the CLIP-ViT method, after PCA is incorporated, outperforms other adapter methods based on the CLIP model. The PCACLIP-3D model achieves the best results of 82.28%.
QV-Pipe. The results of QV-Pipe dataset are reported in Table 2. We can see that our proposed PCACLIP-ViT methods achieve SOTA performance compared to previous methods. It outperforms the ST-Adapter method by +2.0%. Additionally, we also showcase the results of other baseline methods[uniformer, clip, timesformer, r2p1d] with PCA pipeline. QV-Pipe is a complex multilabel dataset requiring global information for perception[videopipe]. The pure convolution network can not generalize well, because convolution has limited capabilities to integrate global information[uniformer]. Therefore, the performance of R(2+1)D[r2p1d] is not as good as transformer-based models with global attention to QV-Pipe. By incorporating the PCA method, our model outperformed the respective baseline models, achieving state-of-the-art performance.
Table 3: Comparison with the state-of-the-art on ARID[arid]
@@
  ST-Adapter[stadapter] 72.67% 97.51%
CLIP-ViT[clip] 77.96% 97.08%
Timesformer[timesformer] 78.53% 96.02%
CLIP-3D[clip] 78.79% 98.44%
EVL[frozenclip] 79.01% 98.44%
UniFormer[uniformer] 86.60% 96.46%
DarkLight-ResNeXt[darklight] 87.27% 99.47%
R(2+1)D[darklight] 90.45% 98.11%
R(2+1)D-Light[darklight] 91.55% 99.51%
R(2+1)D-DarkLight-SA[darklight] 94.04% 99.87%
@@
ARID. We compared our approach with previous mainstream models in the ARID dataset in Table 3. After incorporating the PCA method, the performance of all models surpassed their corresponding baseline models. With the help of text and visual knowledge, PCAR(2+1)D model achieves the best Top-1 Acc of 98.9%. Note that, ARID has the limited number of training videos (less than 4k). In general, ViTs have more training parameters than CNNs, e.g., ViT-Base has 86M model parameters, while R(2+1)D-18 has 33.3M model parameters. Hence, ViTs is more inclined to be overfitting than CNNs, given such limited training data in ARID. As a result, R(2+1)D achieves a better performance than CLIP-ViT. Additionally, ARID refers to dark-illumination video recognition, where action labels are fine-grained with high confusion. This makes CLIP-ViT even worse on Top-1 Acc. Hence, the performance gap between R(2+1)D and CLIP-ViT in Top-1 Acc is larger than that in Top-5 Acc.
@@ 4.2 Ablation Studies
UniFormer[uniformer] $\times$ $\times$ 77.80% 86.60% 96.46% 55.30%
@@
Timesformer[timesformer] $\times$ $\times$ 77.40% 78.53% 96.02% 60.55%
@@
CLIP-ViT[clip] $\times$ $\times$ 74.70% 77.96% 97.08% 63.40%
@@
CLIP-3D[clip] $\times$ $\times$ 78.80% 78.79% 97.44% 55.80%
@@
Multimodal Ablation for PCA: We insert one block of the adapt module to the baseline models for multimodal ablation. The ablation of the block number of the Adapt module is described in detail in the following subsection. According to Table 4, all models have gained performance improvement when incorporating only single-modal knowledge(either visual or text), compared to their corresponding baselines. This indicates that external knowledge, whether in the form of text or visual modalities, has a positive impact on improving the accuracy of the model. Moreover, the integration of both visual and textual knowledge further boosts the models’ performance, showing the effectiveness of our proposed method.
@@
The Number of Inserted Adapt Blocks: To validate the ability of the adapt module to incorporate intermediate-level features into the backbone network, we extracted the intermediate layers of the visual feature extraction models and integrated them into the backbone. Specifically, the block num represents the total number of adapt modules inserted as shown in Fig.4. For the CLIP-3D, Timesformer, and CLIP-ViT models, we extracted the Residual Attention Blocks[vit] features from the middle layers. For the UniFormer model, we extracted middle layer features from SABlock layers[uniformer]. For the R(2+1)D model, we extracted intermediate features from BasicBlocks[r2p1d]. We only introduced textual knowledge in the last block of the model. The results shown in Table 7, Block Num 3 achieves the best result on three datasets. With the increased number of inserted blocks, the results improve.
Results of Visual Enhancement: The first row refers to training the baseline model using the source data. The second row represents training the baseline model using augmented visual enhanced data. The final row demonstrates the effect of incorporating multimodal knowledge into PCA. As shown in Table 8, introducing visual single-modal knowledge in PCA gets better performance, since the PCA visual incorporates information from both the source data and the visual enhanced data. Finally, incorporating knowledge allows the network to achieve superior results.
@@
Table 9: Results of different thresholds in the Chat process
  Threshold TinyVIRAT ARID QV-Pipe
F1-Score Top-1 Acc mAP
  0.2 81.65 98.02 64.55
0.5 82.28 98.90 65.42
0.7 81.21 98.25 64.06
Threshold in the Chat Stage: As its role refers to evaluate the prediction score (i.e., label probability), we choose three thresholds from [0,1] for ablation in Table. 9. We can see that, the performance is the best with the threshold of 0.5. With smaller threshold (e.g., 0.2), the Chat stage tends to use the prompt of the predicted labels as external textual knowledge. With smaller threshold (e.g., 0.7), the Chat stage tends to use the caption of the original video as external textual knowledge. Both cases would introduce unnecessary error, due to the poor confidence of label predictions or the unsatisfactory caption of noisy videos. Hence, to balance prompt knowledge and caption knowledge, the final threshold is chosen as 0.5 for the best result.
[…]
@@ 4.3 Visualization
We visualize the attention regions of the model using Grad-CAM[cam] as shown in Fig.6. After incorporating multi-model knowledge, the model gains the ability to correct the wrong answers provided by the baseline model.
Figure 6: We use CAM[cam] to visualize the PCA pipeline. For each case, the top left corner of each example shows four frames sampled from the original video, while the bottom left corner shows the features obtained without incorporating external knowledge. The images in the middle represents enhanced video frames. The text box contains the textual knowledge obtained by inputting the original video into the VideoChat foundation model. The upper right corner displays features that only include external visual knowledge, while the bottom right corner shows features that only include external textual knowledge. The middle section represents features obtained by incorporating both visual and textual knowledge.
(A) is from TinyVIRAT dataset. When there is no external knowledge, the model predicts the result as ”Opening”. After incorporating visual knowledge or textual knowledge alone, the model also can not predict the correct category. When the external knowledge includes both visual and textual modalities, the focus of the feature maps shifts to the body and feet and predict the correct category.
(B) is from the ARID dataset. baseline model makes incorrect judgments. The enhanced video restores hand movements in a dark environment. Textual knowledge mentions the person using a phone, reading a book, or typing, indicating hand movements. With the incorporation of single-modal knowledge, the model still fails to determine the correct category. After incorporating multimodal knowledge, the model can infer comprehensively that the person is picking something and predict the right category. Furthermore, the feature maps intuitively focus on the position of the hand.
(C) is from the QV-Pipe dataset. Without external knowledge and with only single-modal knowledge, the model cannot predict all class labels. Textual information indicates that this is a sewer with water and some dirt sediment. The model successfully obtains the correct result under the guidance of the semantic segmentation mask and the textual information.
Figure 7: Visualization of the failure cases.
Additionally, we visualize the failure cases of three datasets as shown in Figure.7. The common problem in these cases is that, the ground truth labels are highly confused with other labels in the datasets. Hence, visual and textual knowledge may be insufficient for these challenging cases. It would be interesting to explore other knowledge (like motion, audio, depth) to further enhance our PCA in the future.
@@ 5 Conclusion
In this work, we propose a general knowledge transfer framework PCA, which utilizes a series of foundation models as knowledge containers and progressively incorporates and integrates external multimodal knowledge to enhance the performance of open-world video recognition. We enhance open-world videos to reduce domain gaps and extract external visual knowledge as Percept process. Then, in the Chat stage, we generate rich semantic information as external textual knowledge. Finally, we flexibly fuse external multimodal knowledge into the backbone for open-world recognition. Experimental results demonstrate the effectiveness of our PCA pipeline in open-world scenarios with low-resolution, bad illumination conditions, unusual video scenes, etc. We have not explored how to combine this pipeline for more multi-modal tasks like detection, segmentation, grounding, etc. In future research, it could be explored how this method benefits more multi-modal tasks for open-world video recognition in scenarios like surveillance and robotics.